You want to buy a new vehicle that costs $35,000. You have the option of taking a $3,000 cash back and financing the remaining $32,000 with the dealership for 6% (Option A). The other option is to pay the full $35,000 for the vehicle with an annual financing charge of 1.5% (Option B). Both options allow repayment over 5 years on a monthly basis. Identify the total paid for the car (Option A and B) and the total interest that will be paid for each option (Option A and B) for a 5 year repayment plan. (Round to the nearest dollar 15 not 15.43). There are four questions to answer: Cost of Car option A, Interest on car Option A, Cost of Car Option B, Interest on Car Option B.
Answer #1

Cost of Car Option A=32000*(6%/12)/(1-1/(1+6%/12)^(12*5))*12*5=37118.98

Interest of Car Option A=37118.98-32000=5118.98

Cost of Car Option B=35000*(1.5%/12)/(1-1/(1+1.5%/12)^(12*5))*12*5=36350.76

Interest of Car Option A=36350.76-35000=1350.76